The California Transit system would like to see if the type of transportation used depended on the time of day: Please use the following table to answer questions:

Transportation Total

Bus

Carpool

Own Car

Total

Time of Day

Observed

Expected

Observed

Expected

Observed

Expected

Morning

60

30

30

120

Night

20

20

40

80

Total

80

50

70

200

  1. What is the value of χ2, for independent samples?
  2. What is the critical value of χ2?
  3. What is the conclusion in APA format?
Answer #1

Solution:

Here, we have to use chi square test for independence of two categorical variables.

Null hypothesis: H0: The type of transportation used not depended on the time of day.

Alternative hypothesis: Ha: The type of transportation used depended on the time of day.

We assume level of significance = α = 0.05

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

What is the critical value of χ2?

We are given

Number of rows = r = 2

Number of columns = c = 3

Degrees of freedom = df = (r – 1)*(c – 1) = 1*2 = 2

α = 0.05

Critical value = 5.991465

(by using Chi square table or excel)

Calculation tables for test statistic are given as below:

Observed Frequencies

Type of transportation

Time of Day

Bus

Carpool

Own Car

Total

Morning

60

30

30

120

Night

20

20

40

80

Total

80

50

70

200

Expected Frequencies

Type of transportation

Time of Day

Bus

Carpool

Own Car

Total

Morning

48

30

42

120

Night

32

20

28

80

Total

80

50

70

200

Transportation Total

Bus

Carpool

Own Car

Total

Time of Day

Observed

Expected

Observed

Expected

Observed

Expected

Morning

60

80*120/200 = 48

30

50*120/200 = 30

30

70*120/200 = 42

120

Night

20

80*80/200 = 32

20

50*80/200 = 20

40

70*80/200 = 28

80

Total

80

80

50

50

70

70

200

The variable time of the day represents rows.

The variable type of transportation represents columns.

Calculations

(O - E)

12

0

-12

-12

0

12

(O - E)^2/E

3

0

3.428571

4.5

0

5.142857

What is the value of χ2, for independent samples?

Chi square = ∑[(O – E)^2/E] = 16.07143

P-value = 0.000324

(By using Chi square table or excel)

P-value < α = 0.05

So, we reject the null hypothesis

Conclusion:

There is sufficient evidence to conclude that the type of transportation used depended on the time of day.